OpenStudy (anonymous):

Suppose a factory can have no more than 200 workers on a shift, but must have at least 100 and must manufacture at least 3000 units at minimum cost. The managers need to know how many workers should be on a shift in order to produce the required units at minimal cost. Let “x” represent the number of workers and y represent the number of units manufactured. If the cost per worker is $50 per day and the cost to manufacture 1 unit is $100, write an expression representing the total daily cost C.

OpenStudy (pfenn1):

Let x represent the number of workers and y represent the number of units manufactured. We know from the first statement above that \[100 \le x \le 200\]and \[y \ge 3000\]Total daily cost C will be the daily cost per worker times the number of workers and the cost to manufacture 1 unit times the number of units. Can you write the equation for that?